The are of a rectangle is 45cm square. The shorter sides are 5 cm long, how much longer are the longer sides

Answer by checkley77(12844)

AREA=L*W
L=5
45=5*W
W=45/5
W=9 CM FOR THE WIDTH.
PROOF:
45=5*9
45=45